OpenStudy (anonymous):

Find the equation of the line that passes through the point P(-4; 3) and is parallel to the line y = -3/5x +3. A)y-3=-3/5(X+4) B) y+3=-3/5 (x+4) C)y+3=5/3(x-4) D)y-3=5/3 (x-4)

OpenStudy (anonymous):

when two lines are parallel their slopes are equal first determine the slope of the given line from this slope of the required line will be found then use the slope point form y-y1 = m (x-x1) where m= slope